OpenStudy (anonymous):

the local radio station is giving away some tickets to the circus.adult tickets cost the station $8 each while children's tickets cost $6.the station wants to spend no more than $172.let x equal the number of adults tickets and y equal the number of children's tickets.show all possible solution by graphing

OpenStudy (anonymous):

8x+6y=172 6y=172 - 8x y= -(4/3)x + (172/6) (we divided by 6) Now graph it on a graphing calculator or by hand.

OpenStudy (phi):

I would tweak the answer to y ≤ (-4/3)x + 86/3 as this is an inequality.

OpenStudy (anonymous):

@phi Yeah sorry I didn't the "more than" part